# Iso-Diffusion: Improving Diffusion Probabilistic Models Using the Isotropy of the Additive Gaussian Noise

### v1

@@ Abstract
Denoising Diffusion Probabilistic Models (DDPMs) have accomplished much in the realm of generative AI. Despite their high performance, there is room for improvement, especially in terms of sample fidelity by utilizing statistical properties that impose structural integrity, such as isotropy. Minimizing the mean squared error between the additive and predicted noise alone does not impose constraints on the predicted noise to be isotropic. Thus, we were motivated to utilize the isotropy of the additive noise as a constraint on the objective function to enhance the fidelity of DDPMs. Our approach is simple and can be applied to any DDPM variant. We validate our approach by presenting experiments conducted on four synthetic 2D datasets as well as on unconditional image generation. As demonstrated by the results, the incorporation of this constraint improves the fidelity metrics, Precision and Density for the 2D datasets as well as for the unconditional image generation.
Keywords:
Generative Models Diffusion Models Structure Image Generation Gaussian Noise
@@ 1 Introduction
(a) DDPM
(b) Ours $\lambda=0.01$
(c) Ours $\lambda=0.05$
(d) Ours $\lambda=0.1$
(e) Ours $\lambda=0.2$
(f) Ours $\lambda=0.3$
Figure 1: The variation of the Generated Distribution (in blue) compared to the Ground Truth Distribution (in orange). The plots indicate that the generated samples are much more closely connected and densely packed with the increase of the regularization parameter, $\lambda$. The plots show the utility of imposing the isotropy constraint based loss function which enables the generation of more realistic samples as most of generated samples are concentrated near the ground truth samples.
@@
Diffusion models have been accomplishing great feats in the realm of generative AI, specifically in terms of unconditional and conditional image generation ([12], [5], [18], [11], [15], [17], [6]). Starting with the revolutionary paper by Ho et al. [4] and the improvements by Nichol et al. [12] as well as the Latent Diffusion Model by Rombach et al. [17], these models have had the biggest impact in this context. The fidelity and diversity of these generated images are surprisingly amazing. Yet, as with all models, these models can still be improved upon closer inspection. As with the improvements done by Nichol et al. [12] to the original Denoising Diffusion Probabilistic Model (DDPM) by introducing techniques such as the cosine-based variance schedule as well as making the model learn the variance rather than keeping it fixed, helped improve the performance of DDPMs. Our goal in this paper is to make a similar contribution with regards to the improvement of the important fidelity metrics, Density [10] and Precision [8].
@@
Although DDPMs perform well, we noticed that these existing models do not necessarily incorporate any distributional (structural) information about the particular dataset it tries to generate. Typically, the DDPM’s forward process gradually pushes the dataset towards a white gaussian, which can be thought of as a structural vanishing point of the data distribution. This implies a well placed point of origin for the generative process (reverse path) from a point of complete lack of structure towards the final destination which is the data distribution. In the DDPM implementation, the learning process considers the expected squared norm difference between the additive gaussian noise and the predicted noise as its objective function. Therefore, for the generative process, to enhance the aforementioned creation of structure, the objective function can be modified to include any structural measure, such as isotropy. In the context of this paper, "isotropy" describes the affinity towards a white distribution quantified by the expected squared norm of the distribution. However, the objective function in the existing models only focus on the expected squared norm difference between the additive gaussian noise and the predicted noise.
@@
Thus, we were motivated to include the isotropic nature of the additive gaussian noise when optimizing for the objective to further enhance the statistical properties of the predicted noise. Our intuition is that by capturing the statistical properties of the noise in more detail, the model will be able to produce higher fidelity samples as it would have much more information regarding the distributional structure of the samples.
@@
As the rationale for introducing isotropy to the objective function has been established. Now, we let’s see how isotropy establishes convergence and quantifies structural information about the distribution. For example, the isotropy of a white gaussian vector in $\mathbb{R}^{n}$ is the expected squared norm of that vector, which is equal to its dimension, $n$ [23]. This establishes the upper bound in the limit for a normalized distribution with complete lack of structure, which in other words is white. On the hand, the desired distribution, which has more structure and is more colored, would consequently have a lower isotropy value. This implies that the generative process, in its drive towards a structural distribution minimizes isotropy. Furthermore, when analyzing the mean square error objective, we observed that the isotropic nature of the noise, when included, effectively makes the objective function equal to 0 in expectation.
@@
The inclusion of this constraint does not incur a large computational cost and can be readily applied to any of the diffusion model variants. In this work, we experimented on four 2D synthetic datasets with our modified objective function and show that the fidelity metrics, in particular the Density, improves significantly. This implies that the proposed constraint attempt to latch on to the more information rich dense modes of the desired distribution. Furthermore, we validate our approach on unconditional image generation using the Oxford Flowers [13] and Oxford-IIIT Pet [14] datasets. We compare the fidelity and diversity of the generated samples based on key evaluation metrics such as Precision and Recall [8], Density and Coverage [10], Frechet Inception Distance (FID) [2] and Inception Score (IS) [19].
@@
  • We introduce Iso-Diffusion: a modified approach that introduces an isotropic constraint on the predicted noise objective function to steer the generative process in a structurally coherent manner. This results in improved fidelity of the generated data distribution. We believe, to the best of our knowledge, that we are the first to propose such a modified loss based on the properties of the noise.
  • We analyze the simple loss function in the DDPM and its connection to isotropy. Moreover, we show that the isotropy of the data distribution monotonically increases and converges to the maximum isotropy value which corresponds to a white gaussian distribution. This confirms that the definition of isotropy, mentioned in this paper, conveys information about the structural information of the data distribution when the data distribution undergoes the forward process in DDPMs.
  • We evaluate and validate our approach on four 2D synthetic datasets as well as on the task of unconditional image generation on Oxford Flowers and Oxford-IIIT Pet datasets. Considering the key evaluation metrics, such as the Precision, Recall, Density, Coverage, FID and IS, the modified objective is able to surpass the original DDPM with a significant gap in terms of the fidelity metrics, Density and Precision.
@@ 2 Related Work
Deep Generative Models Generative models (GANs [1], VAEs [7], flow-based models [16], and diffusion models [4]) learn the probability distribution of given data, allowing us to sample new data points from the distribution. Deep generative models have been used for generating images, videos [3], 3d objects [9], etc. Moreover, these models have been used for inverse problem solving [22] and to understanding the latent representations of the distributions.
@@
However, most of these improvements were focused on improving the models based on the most widely used metrics for image generation, FID and IS. But, some of the recent work ([8], [10]), in generative models have pointed out that FID and IS are not necessarily indicative of the actual fidelity of the samples generated by generative models. Thus, researchers have been focusing on finding other metrics, such as Precision and Density, to assess the fidelity of these generated samples. In particular we observed that the Density takes the local context (measuring how close it is to densely packed samples of the true distribution) of a sample into account during its calculation. We believe that this makes the Density a vital metric to asses the samples’ fidelity.
@@ 3 Background
Diffusion probabilistic models were first introduced by Sohl-Dickstein et al. [20] These models fall under the category of generative models which learn the distribution of the data so that they can sample from these data distributions. However, it wasn’t until Ho et al. [4] that Diffusion Probabilistic Models took off. In the next few subsections, we will provide a brief overview of the DDPM definitions that will be useful to understanding our work.
@@ 3.1 Definitions
In the DDPM, we simply add a gaussian noise, which varies according to a specific variance schedule, $\beta_{t}\in(0,1)$. The noise at each time-step corrupts the data, such that by the time the time-step reaches its final value, $T$, the data will be mapped to an almost white gaussian distribution. However, the learning occurs when we try to learn the reverse process by which we try to denoise along the same trajectory starting from the almost white gaussian distribution. The first process, in which we add noise, is called the forward process and the latter, in which we denoise, the reverse process. The forward process is often characterized by $q$ and the reverse process by $p$. Both of which are modeled as gaussian distributions.
@@
The reverse process, given by $p\sim\mathcal{N}(x_{t-1}|x_{t})$, can be obtained in terms of the forward process distribution $q$ and the Baye’s Theorem. However, the reverse process only becomes tractable when the posterior distribution $q(x_{t-1}|x_{t})$, is conditioned on the input data $x_{0}$. Thus, during training, the model tries to learn the tractable $q(x_{t-1}|x_{t},x_{0})$ distribution. This distribution, which is also a gaussian distribution, is defined by the following equation and parameters.
@@ 3.2 Training Process
To train, however, one could make the model predict the mean of the reverse process distribution at each time step. But, Ho et al. [4] mentions that predicting, the additive noise, $\epsilon$, leads to better results. The additive noise and the mean of the reverse process distribution at each time step are elegantly linked by equations (5) and (8). This results in the following re-parametrization of $\tilde{\mu}(x_{t},t)$,
@@
Therefore, predicting the additive noise, $\epsilon$, is adequate for the task of predicting the mean of the backward process distribution. Moreover, since the forward process’ variance schedule is fixed, the reverse process variance, $\tilde{\beta}_{t}$, is also assumed to be fixed according to $\tilde{\beta}_{t}$.
@@
where, $\epsilon_{\theta}(x_{t},t)$ is the predicted noise.
@@ 3.3 Hidden Statistical Properties of $\epsilon$
Furthermore, since the goal is to predict the noise as accurately as possible, $\epsilon_{\theta}$ should also be distributed according to a white gaussian distribution, i.e., $\epsilon_{\theta}\sim\mathcal{N}(0,\mathbf{I})$). Hence, if $\epsilon$ and $\epsilon_{\theta}$ are both independent identical isotropic random vectors,
@@ 4 Analysis on the Isotropy of $x_{t}$
Inspired, we wanted to find out further implications of imposing structural information in the DDPM. As it turns out, we were able to gain more interesting insights about the forward process of the DDPM. For example, if we consider equation (5) and consider the isotropy, expected squared norm of $x_{t}$, we see that,
$E(||x_{t}||^{2})=E(x_{t}^{T}x_{t})\\ =\bar{\alpha}_{t}E(x_{0}^{T}x_{0})+(1-\bar{\alpha}_{t})E(\epsilon^{T}\epsilon)+2\sqrt{(\bar{\alpha}_{t})(1-\bar{\alpha}_{t})}E(x_{0}^{T}\epsilon)$ (16)
However, since $\epsilon$ is a white gaussian random vector, it is isotropic. Moreover, by assuming that it is independent of the distribution of $x_{0}$, when $x_{0}$ is non-isotropic, we see that,
@@
Therefore, for any particular normalized data distribution, we see that during the forward process, the isotropy of the data distribution increases, and finally converges to the isotropy of a white gaussian vector, when the data distribution completely converts into a white gaussian distribution. Hence, the definition of isotropy given in this paper, aligns perfectly with the fact that the isotropy quantifies structural information about the data distribution.
@@ 5 Isotropy Based Loss Function
Armed with the above analyses, we proceeded to modify the objective function $L_{\text{simple}}$ to include a regularization term which penalizes the model, if the model predicts a noise which is not necessarily isotropic. Hence, the new modified objective function we propose to optimize is,
@@
However, this modified objective needs to be further simplified so as to make this new error be independent of the size of the dimension of the random vector. Thus, we make the following modification during implementation.

### v2

@@ Abstract
Denoising Diffusion Probabilistic Models (DDPMs) have accomplished much in the realm of generative AI. With the tremendous level of popularity the Generative AI algorithms have achieved, the demand for higher levels of performance continues to increase. Under this backdrop, careful scrutinization of algorithm performance under sample fidelity type measures is essential to ascertain how, effectively, the underlying structures of the data distribution were learned. In this context, minimizing the mean squared error between the additive and predicted noise alone does not impose structural integrity constraints on the predicted noise, for instance, isotropic. Under this premise, we were motivated to utilize the isotropy of the additive noise as a constraint on the objective function to enhance the fidelity of DDPMs. Our approach is simple and can be applied to any DDPM variant. We validate our approach by presenting experiments conducted on four synthetic 2D datasets as well as on unconditional image generation. As demonstrated by the results, the incorporation of this constraint improves the fidelity metrics, Precision and Density, and the results clearly indicate how the structural imposition was effective.
@@ 1 Introduction
Diffusion models have been accomplishing great feats in the realm of generative AI, specifically in terms of unconditional and conditional image generation ([19], [9], [26], [18], [22], [24], [6], [10], [1]). Starting with the revolutionary paper by Ho et al. [7] and the improvements by Nichol et al. [19] as well as the Latent Diffusion Model by Rombach et al. [24], these models have had the biggest impact in this context. The fidelity and diversity of the images generated by these models are surprisingly amazing. Yet, as with all models, these models can still be improved upon closer inspection. As with the improvements done by Nichol et al. [19] to the original Denoising Diffusion Probabilistic Model (DDPM) by introducing techniques such as the cosine-based variance schedule and allowing the model to learn the variance rather than keeping it fixed helped improve the performance of DDPMs. Our goal in this paper is to make a similar contribution with regard to the improvement of the important fidelity metrics, Density [16] and Precision [13], by imposing possible regularizations that promote the modified DDPM algorithm to learn the underlying structures, diversity, modality and density spread of the true distribution.
@@
Although DDPMs perform well, we noticed that these existing models do not necessarily incorporate any distributional (structural) information about the particular dataset it tries to generate. Typically, the DDPM’s forward process gradually pushes the dataset towards an isotropic Gaussian, which can be thought of as the structural vanishing point of the data distribution [17]. This implies a well-placed point of origin for the generative process (reverse path) from a point of complete lack of structure toward the final destination, which is the data distribution. In the DDPM implementation, the learning process considers the expected squared norm difference between the additive Gaussian noise and the predicted noise as its objective function. Therefore, for the generative process, to enhance the aforementioned creation of structure, the objective function can be modified to include any structural measure, such as isotropy.
@@
Thus, we were motivated to include the isotropic nature of the additive Gaussian noise when optimizing for the objective to further enhance the statistical properties of the predicted noise through backpropagation. The current objective function of the DDPM does not include any mechanism that explicitly encourages the isotropic nature of the predicted noise. Therefore, a mechanism that guarantees the model progresses from a more non-isotropic distribution (distributions with multiple modes, discontinuities or non-uniformities of density distributions, non-uniformly distributed spatial structures) to an isotropic Gaussian distribution toward the vanishing point in a structured and learned manner is needed. Our intuition is that by capturing the statistical properties of the noise in more detail, the model will be able to produce higher-fidelity samples as it would have much more information regarding the distributional structure of the samples.
@@
As the rationale for introducing isotropy to the objective function has been established, now, let us see how isotropy establishes convergence and quantifies structural information about the distribution. For example, the isotropy of an isotropic random vector in $\mathbb{R}^{n}$ is the expected squared norm of that vector, which is equal to its dimension, $n$ [34]. This establishes the convergence in the limit for a normalized distribution with a complete lack of structure, which in other words is isotropic. Conversely, the desired distribution, which has more structure and is more non-isotropic, would consequently have a lower isotropy value. This implies that the generative process, in its drive towards a structural distribution, minimizes isotropy. Furthermore, when analyzing the mean square error objective, we observed that incorporating the isotropic nature of the noise effectively makes the objective function equal to zero in expectation.
@@
The inclusion of this constraint does not incur a large computational cost and can be readily applied to any of the diffusion model variants. In this work, we scrutinized the DDPM model’s behavioral aspects to interpret its functionality using well-defined 2D synthetic datasets, such as Swiss Roll, Scattered Moon, Moon with two circles and Central Banana, to draw fundamental conclusions about DDPM algorithm. Furthermore, we experimented on four 2D synthetic datasets with our modified objective function and showed that the fidelity metrics, in particular the Precision and Density, improved significantly. In addition, we validate our approach to unconditional image generation using the Oxford Flower [20] and Oxford-IIIT Pet [21], CIFAR-10 [12] and CIFAR-100 [12] datasets. We compare the fidelity and diversity of the generated samples based on key evaluation metrics such as Precision and Recall [13], Density and Coverage [16], Frechet Inception Distance (FID) [5] and Inception Score (IS) [28].
@@
  • We introduce Iso-Diffusion: a modified approach that introduces an isotropic constraint on the predicted noise objective function to steer the generative process in a structurally coherent manner. This results in improved fidelity of the generated data distribution. We believe, to the best of our knowledge, that we are the first to propose such a modified loss based on the structural properties of the noise.
  • We analyze the simple loss function in the DDPM and its connection to isotropy. Moreover, we show that the isotropy of the data distribution monotonically increases and converges to the maximum isotropy value, which corresponds to an isotropic Gaussian distribution. This confirms that the definition of isotropy mentioned in this paper, conveys information about the structure of the data distribution when the data distribution undergoes the forward process in DDPMs.
  • We evaluate and validate our approach on four 2D synthetic datasets as well as on the task of unconditional image generation on Oxford Flower, Oxford-IIIT Pet, CIFAR-10 and CIFAR-100 datasets. Considering the key evaluation metrics, such as Precision, Recall, Density, Coverage, FID and IS, the modified objective is able to surpass the original DDPM with a significant gap in terms of the fidelity metrics, Density and Precision.
  • We conduct an in-depth analysis of the Density and Coverage metrics to evaluate the generative capabilities of Iso-Diffusion compared to DDPM. This analysis facilitates a detailed comparison between the generated and true data distributions, visually illustrating Iso-Diffusion’s superior alignment with the true distribution. Furthermore, it highlights the importance of these metrics for assessing generative AI algorithms in computer vision applications.
@@ 2 Related Work
Generative models, particularly in recent years, have gained significant momentum due to their applications in various fields. They began with specific use cases and have evolved along a clear trajectory, as outlined below.
Deep Generative Models Generative models (GANs [3], VAEs [11], flow-based models [23], and diffusion models [7]) learn the probability distribution of given data, allowing us to sample new data points from the distribution. Deep generative models have been used for generating images, videos [8], 3d objects [15], etc. Moreover, these models have been used for inverse problem solving [33] [14] and to understanding the latent representations of the distributions.
@@
However, most of these improvements were focused on improving the models based on the most widely used metrics for image generation, FID and IS. But some of the recent work ([13], [16], [25]), in generative models has pointed out that FID and IS are not necessarily indicative of the actual fidelity of the samples generated by generative models. Thus, researchers have been focusing on finding other metrics, such as Precision and Density, to assess the fidelity of these generated samples [2], [30]. In particular, we observed that the Density takes the local context (measuring how close it is to densely packed samples of the true distribution) of a sample into account during its calculation. We believe that this makes the Density a vital metric to assess the samples’ fidelity.
@@ 3 Background
Figure 2: Variation of isotropy of the data distribution along the forward diffusion process for the 2D synthetic datasets. As can be seen from the plot, in the limit, the data distribution reaches the value of two, which happens to be the dimension of an isotropic random vector in $\mathbb{R}^{2}$ (expected squared norm of an isotropic random vector).
Figure 3: An example scenario for illustrating a situation where high Density and low Coverage is recorded. Generating samples in the neighborhoods of the highly dense regions over the outliers in the true manifold has resulted in a high Density and low Coverage.
Diffusion probabilistic models were first introduced by Sohl-Dickstein et al. [31] These models fall under the category of generative models which learn the distribution of the data so that they can sample from these data distributions. However, it was not until Ho et al. [7] that Diffusion Probabilistic Models took off. In the next few subsections, we will provide a brief overview of the DDPM definitions that will be useful to understanding our work.
@@ 3.1 Definitions
In the DDPM, we simply add a Gaussian noise, which varies according to a specific variance schedule, $\beta_{t}\in(0,1)$. The noise at each time-step corrupts the data, such that by the time the time-step reaches its final value, $T$, the data will be mapped to an almost isotropic Gaussian distribution. However, the learning occurs when we try to learn the reverse process by which we try to denoise along the same trajectory starting from the almost isotropic Gaussian distribution. The first process, in which we add noise, is called the forward process and the latter, in which we denoise, is called the reverse process. The forward process is often characterized by $q$ and the reverse process by $p$. Both of which are modeled as Gaussian distributions.
@@
The reverse process, given by $p\sim\mathcal{N}(x_{t-1}|x_{t})$, can be obtained in terms of the forward process distribution $q$ and Baye’s Theorem. However, the reverse process only becomes tractable when the posterior distribution $q(x_{t-1}|x_{t})$, is conditioned on the input data $x_{0}$. Thus, during training, the model tries to learn the tractable $q(x_{t-1}|x_{t},x_{0})$ distribution. This distribution, which is also a Gaussian distribution, is defined by the following equation and parameters.
@@ 3.2 Training Process
To train, however, one could make the model predict the mean of the reverse process distribution at each time step. But Ho et al. [7] mentions that predicting the additive noise, $\epsilon$, leads to better results. The additive noise and the mean of the reverse process distribution at each time step are elegantly linked by equations 5 and 8. This results in the following re-parametrization of $\tilde{\mu}(x_{t},t)$,
@@
Therefore, predicting the additive noise $\epsilon$, is adequate for the task of predicting the mean of the backward process distribution. Moreover, since the forward process’ variance schedule is fixed, the reverse process variance, $\tilde{\beta}_{t}$, is also assumed to be fixed according to $\tilde{\beta}_{t}$.
@@
where $\epsilon_{\theta}(x_{t},t)$ is the predicted noise.
@@ 3.3 Hidden Statistical Properties of $\epsilon$
As discussed earlier, one of the main objectives of the proposed algorithm is to identify a learnable isotropic measure that best reflects the overall isotropic nature of the learned samples. This will allow backpropagation to accurately guide the model toward the maximum isotropy at the vanishing point. The identified metric (11) is the expected squared norm of $\epsilon$ which will be named isotropy.
$\textit{Isotropy}=E(\epsilon^{T}\epsilon)$ (11)
@@
Furthermore, since the goal is to predict the noise as accurately as possible, $\epsilon_{\theta}$ should also be distributed according to an isotropic Gaussian distribution, i.e., $\epsilon_{\theta}\sim\mathcal{N}(0,\mathbf{I})$. Hence, if $\epsilon$ and $\epsilon_{\theta}$ are both identical isotropic random vectors,
@@ 4 Analysis on the Isotropy of $x_{t}$
Based on the observations, we were inspired to find out further implications of imposing structural information in the DDPM. As it turns out, we were able to gain more interesting insights about the forward process of the DDPM. For example, if we consider equation 5 and consider the isotropy, expected squared norm of $x_{t}$, we see that,
$E(||x_{t}||^{2})$ $=E(x_{t}^{T}x_{t})$
$=\bar{\alpha}_{t}E(x_{0}^{T}x_{0})+(1-\bar{\alpha}_{t})E(\epsilon^{T}\epsilon)$
$\quad+2\sqrt{(\bar{\alpha}_{t})(1-\bar{\alpha}_{t})}E(x_{0}^{T}\epsilon)$ (17)
However, since $\epsilon$ is an isotropic Gaussian random vector, it is isotropic. Moreover, by assuming that it is independent of the distribution of $x_{0}$, when $x_{0}$ is non-isotropic, we see that,
@@
Therefore, for any particular normalized data distribution, we see that during the forward process, the isotropy of the data distribution increases. Finally, converges to the isotropy of an isotropic random Gaussian vector, when the data distribution completely converts into an isotropic Gaussian distribution (see Figure 2). Hence, the definition of isotropy given in this paper aligns perfectly with the fact that the isotropy quantifies structural information about the data distribution.
@@ 5 Isotropy Based Loss Function
Metrics Swiss Roll Scattered Moon Moon with two circles Central Banana
DDPM Ours DDPM Ours DDPM Ours DDPM Ours
Precision 0.9458 0.9893 (+4.60%) 0.9990 0.9993 (+0.03%) 0.9921 0.9982 (+0.61%) 0.8974 0.9072 (+1.09%)
Recall 0.9927 0.9709 (-2.19%) 0.9962 0.9736 (-2.27%) 0.9967 0.9694 (-2.74%) 0.9977 0.9417 (-5.61%)
Density 0.8946 0.9908 (+10.75%) 1.0015 1.0049 (+0.34%) 0.9925 1.0081 (+1.57%) 0.8785 0.8962 (+2.01%)
Coverage 0.8932 0.8458 (-5.31%) 0.9605 0.8254 (-14.07%) 0.9498 0.8572 (-9.75%) 0.9102 0.6840 (-24.85%)
Metrics Oxford Flower Oxford-IIIT-Pet CIFAR-10 CIFAR-100
DDPM Ours DDPM Ours DDPM Ours DDPM Ours
FID ($\downarrow$) 55.590 47.310 (-14.9%) 34.087 31.900 (-6.4%) 16.023 11.872 (-25.9%) 14.794 14.141 (-4.4%)
IS ($\uparrow$) 3.097 3.504 (+13.1% 7.083 7.531 (+6.3%) 8.463 8.482 (+0.2%) 9.032 9.183 (+1.7%)
Precision ($\uparrow$) 0.725 0.944 (+30.3%) 0.819 0.954 (+16.5%) 0.607 0.689 (+13.6%) 0.638 0.710 (+11.4%)
Recall ($\uparrow$) 0.184 0.056 (-69.8%) 0.152 0.063 (-58.4%) 0.447 0.384 (-14.0%) 0.398 0.350 (-12.1%)
Density ($\uparrow$) 2.632 11.039 (+319.4%) 6.704 15.778 (+135.4%) 1.401 2.104 (+50.3%) 1.479 2.190 (+48.1%)
Coverage ($\uparrow$) 0.959 0.994 (+3.6%) 0.9996 0.9999 (+0.03%) 0.987 0.995 (+0.8%) 0.9996 0.9996 (0.0%)
Table 2: Comparison of Evaluation Metrics for the two methods: DDPM and Iso-Diffusion for the Image Datasets.
In the default DDPM model, the variance schedule drives the transformation toward an isotropic Gaussian distribution by restricting the degrees of freedom for the movement of information of the distribution, without using backpropagation to adaptively learn the degree of isotropy achieved, making it, a non-learnable process. Armed with the above analyses, we proceeded to modify the objective function $L_{\text{simple}}$ to include a regularization term which penalizes the model, if the model predicts a noise which is not necessarily isotropic. Hence, the new modified objective function we propose to optimize is,
@@
However, this modified objective needs to be further simplified so as to make this new error, independent of the size of the dimension of the random vector. Thus, we make the following modification during implementation.
@@
6 Interpret the Evaluation Metrics
Precision denotes the fraction of generated data that lies in the true manifold, by counting whether each generated data point falls within a neighborhood sphere of real samples. This measure reflects how closely the generated points align with the true manifold [13], [27].
Recall denotes the fraction of true data that lies in the generated manifold, by counting whether each true data point falls within a neighborhood sphere of generated samples. This measure indicates how well the true points align with the generated manifold [13], [27].
Density counts the number of neighborhood spheres of real samples that encompass each generated data point. This allows Density to reward generated samples located in areas densely populated by real samples, reducing sensitivity to outliers. This enables to consider the local context of a distribution by measuring how close a sample is to densely packed points in the true distribution [16].
Coverage measures the fraction of real samples whose neighborhoods contain at least one generated sample. Moreover, Coverage measures the diversity of a distribution by assessing whether all aspects of the distribution are represented. However, the presence of sparse outliers in the true manifold and the absence of the generated samples near those outliers may result in low Coverage [16]. (see Figure 3)